Sisyphus is towing a 60-kilogram ball up a 27° incline with a constant acceleration of 1.3...

  1. Sisyphus is towing a 60-kilogram ball up a 27° incline with a constant acceleration of 1.3 m/s2.  The coefficient of friction between the ball and the slope is 0.33.  Find the tension in the cable that connects Sisyphus to the lead ball.
Answer #1

here,

the mass of lead ball , m = 60 kg

theta = 27 degree

constant acceleration , a = 1.3 m/s^2

the coefficient of friction , uk = 0.33

let the tension in the cable be T

for the ball

T - m * g * sin(theta) - uk * m * g * cos(theta) = m * a

T = 60 * ( 9.81 * sin(27) + 0.33 * 9.81 * cos(27) + 1.3) N

T = 518.3 N

the tension in the cable is 518.3 N